KpopAsianDramaFanYT:

What is the average rate of change from x = 1 to x = 3?

KpopAsianDramaFanYT:

1 attachment
simsharrison:

@jhonyy9

jhonyy9:

do you know what is an average rate of change ? @AZ

AZ:

average rate of change is the slope

AZ:

What class are you taking? @KpopAsianDramaFanYT Are you in calculus or an algebra class?

KpopAsianDramaFanYT:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ What class are you taking? @KpopAsianDramaFanYT Are you in calculus or an algebra class? \(\color{#0cbb34}{\text{End of Quote}}\) Algebra 2

simsharrison:

We learned this in algebra 2, ik my comment probably wasn't necassary

AZ:

so we won't pull out the big boy guns to solve this they just want us to find the slope between (1, -3) and (3, 1)

KpopAsianDramaFanYT:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ average rate of change is the slope \(\color{#0cbb34}{\text{End of Quote}}\) I Kow that but it is confusing

AZ:

\(\text{Slope} = \dfrac{y_2 - y_1}{x_2-x_1}\) where you have two points \(\Large (x_1, y_1)\) and \(\Large (x_2, y_2)\)

KpopAsianDramaFanYT:

I think i got it now

AZ:

I'll throw in colors for you \(\Large\text{Slope} = \dfrac{\color{green}{y_2} - \color{orange}{y_1}}{\color{blue}{x_2}-\color{red}{x_1}}\) where you have two points \(\Large (\color{red}{x_1}, \color{orange}{y_1})\) and \(\Large (\color{blue}{x_2}, \color{green}{y_2})\) So you're two points are \((\color{red}{1}, \color{orange}{-3})\) and \((\color{blue}{3}, \color{green}{1})\)

AZ:

you can use the colors to see where you're supposed to plug in the numbers

KpopAsianDramaFanYT:

THX

AZ:

You're welcome!

KpopAsianDramaFanYT:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KpopAsianDramaFanYT \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ What class are you taking? @KpopAsianDramaFanYT Are you in calculus or an algebra class? \(\color{#0cbb34}{\text{End of Quote}}\) I'm in 8th grade and taking algebra 2

AZ:

Got it. Let me know if you want me to check your answer :)

KpopAsianDramaFanYT:

k

KpopAsianDramaFanYT:

is this correct?: Slope = 1 - (-3) / 3−1 Slope = 4/2 Slope = 2

AZ:

That is absolutely correct!

KpopAsianDramaFanYT:

THX

AZ:

It was my pleasure!